OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 8y = –2 5x + 6y = 9 A. (3, –1) B. (–2, 2) C. (–2, 7) D. (–1, 3)

OpenStudy (anonymous):

@DSS

OpenStudy (anonymous):

@robtobey @Compassionate @campbell_st

OpenStudy (anonymous):

lil help

OpenStudy (anonymous):

You can use elimination. To do this you must get one variable to cancel out. Try this: 2x + 8y = –2 -5(2x+8y=-2) -10x-40y=10 5x + 6y = 9 2(5x+6y=9) 10x + 12y = 18 -28y = 28 y = -1 You now have y...substitute -1 into the equations to get x. 2x+8(-1)=-2 5x+6(-1)=9 2x - 8 = -2 5x-6=9 2x=6 5x=15 x=3 x=3 It works in both equations so (3,-1)

OpenStudy (anonymous):

@DSS thx

OpenStudy (anonymous):

You're welcome